OpenStudy (anonymous):

what is the integral of (sqrt(x))(x+1) over 0,1

OpenStudy (amoodarya):

substitute sqrt(x)=u so x=u^2 dx=2u du apply simplify

OpenStudy (amoodarya):

or distribute sqrt(x)(x+1) =x ^(3/2) +x^1/2 then intergrate